OpenStudy (mark23):

Find the value of x. Round your answer to the nearest tenth. sin 33=7/x

OpenStudy (lapinblanc):

Can you use a calculator for the sin 33 part or are you required to show that work?

OpenStudy (lapinblanc):

Because if you are allowed to use a calculator for the sin value, the answer would be the following: sin33=7/x multiply by reciprocal of x. (x/1)sin33=(7/x)(x/1) this cancels the x on the right and leaves it on the left xsin33=7 multiply by reciprocal of sin33. (1/sin33)(xsin33)=7(1/sin33) sin33 cancels on left and remains in the denominator on the right. Answer: x=7/sin33 Complete and round with calculator.